# EEG-SCMM: Soft Contrastive Masked Modeling for Cross-Corpus EEG-Based Emotion Recognition

### v1

@@ Abstract
Emotion recognition using electroencephalography (EEG) signals has garnered widespread attention in recent years. However, existing studies have struggled to develop a sufficiently generalized model suitable for different datasets without re-training (cross-corpus). This difficulty arises because distribution differences across datasets far exceed the intra-dataset variability. To solve this problem, we propose a novel Soft Contrastive Masked Modeling (SCMM) framework. Inspired by emotional continuity, SCMM integrates soft contrastive learning with a new hybrid masking strategy to effectively mine the ”short-term continuity” characteristics inherent in human emotions. During the self-supervised learning process, soft weights are assigned to sample pairs, enabling adaptive learning of similarity relationships across samples. Furthermore, we introduce an aggregator that weightedly aggregates complementary information from multiple close samples based on pairwise similarities among samples to enhance fine-grained feature representation, which is then used for original sample reconstruction. Extensive experiments on the SEED, SEED-IV and DEAP datasets show that SCMM achieves state-of-the-art (SOTA) performance, outperforming the second-best method by an average accuracy of 4.26% under two types of cross-corpus conditions (same-class and different-class) for EEG-based emotion recognition.
Introduction
@@
Emotions are human attitudinal experiences and behavioral responses to objective things, closely related to health conditions and behavioral patterns (Wang, Zhang, and Tang 2024). Compared to speech (Singh and Goel 2022), gestures (Noroozi et al. 2018), and facial expressions (Canal et al. 2022), electroencephalography (EEG) offers a more direct and objective measurement of human emotions by capturing brain activity across various scalp locations (Hu et al. 2019). Recently, researchers have increasingly emphasised EEG-based emotion recognition (Zhong, Wang, and Miao 2020; Zhao, Yan, and Lu 2021; Zhang, Liu, and Zhong 2022; Gao et al. 2024), aiming to advance the development of affective brain-computer interfaces (aBCIs). However, three critical challenges remain to be addressed in current approaches.
(1) Dataset Specificity. Most existing EEG-based emotion recognition methods are typically designed for a single dataset, necessitating model retraining when the dataset changes. This requirement significantly limits the model’s generalizability and scalability, hindering its application across different datasets. To tackle this issue, the concept of cross-corpus has been proposed, which is designed to be generalized across multiple datasets. A cross-corpus model is trained on one dataset and can be directly applied to another without the need for retraining from scratch. This concept, which originated in natural language processing (Schuller et al. 2010; Zhang et al. 2011), has been extended to various domains in recent years (Rayatdoost and Soleymani 2018; Chien, Yang, and Lee 2020; Ryumina, Dresvyanskiy, and Karpov 2022). Although existing EEG-based emotion recognition methods, such as BiDANN (Li et al. 2018), TANN (Li et al. 2021), and PR-PL (Zhou et al. 2023), have achieved superior performance in within-subject or cross-subject tasks, their efficacy significantly degrades in cross-corpus scenarios, where differences in data distribution across datasets far exceed the variability within a single dataset (Rayatdoost and Soleymani 2018).
@@
Figure 1: (a) An illustration of emotional continuity. We take the sample $x_{i}$ at second $t$ within an EEG trial as the anchor, and calculate the cosine similarity between $x_{i}$ and its neighboring sample $x_{j}$, as well as the distant sample $x_{k}$. High cosine similarities indicate that human emotions remain relatively stable and similar over a certain period. (b) Hard CL vs. SCMM. Traditional hard contrastive learning considers the embeddings $z_{i}$ and $\widetilde{z}_{i}$ of the same sample $x_{i}$ and its augmented view $\widetilde{x}_{i}$ as positive pairs, while different samples and their augmented views are treated as negative pairs. When computing the contrastive loss, the weights for positives and negatives are set to 1 and 0, respectively. In contrast, the soft contrastive learning strategy designed in SCMM generates soft assignments for different sample pairs based on their distances in the original data space.
(2) Data Availability. Current approaches for cross-domain or cross-corpus EEG-based emotion recognition rely heavily on domain adaptation techniques, which depend extensively on the availability of labeled source data and unlabeled target data. For example, AD-TCN (He, Zhong, and Pan 2022) learned an asymmetric mapping that adapts the target domain feature encoder to the source domain, aiming to eliminate the complicated steps of target domain labeling and improve the model performance in cross-domain scenarios. Similarly, E2STN (Zhou et al. 2024) integrated content information from the source domain with style information from the target domain to create stylized emotional EEG representations. However, these methods require prior access to all labeled source data and unlabeled target data for model training, presenting a significant limitation due to data availability constraints (Liu et al. 2024a).
@@
(3) Ignorance of Emotional Continuity. Unlike domain adaptation techniques, contrastive learning (CL) achieves superior performance without relying on labeled data, and has demonstrated significant potential in enhancing feature representation capabilities across various domains (Chen et al. 2020; Radford et al. 2021; Zhang et al. 2022). Current CL-based methods for EEG-based emotion recognition consider an anchor and its augmented views as positive pairs, while treating all other samples as negatives, as shown in Fig. 1(b) (Hard CL). For example, CLISA (Shen et al. 2022) employed contrastive learning to minimize the inter-subject differences by maximizing the similarity in EEG representations across subjects. JCFA (Liu et al. 2024a) performed joint contrastive learning across three domains to align the time- and frequency-based embeddings of the same sample in the latent time-frequency space, achieving state-of-the-art (SOTA) performance in cross-corpus scenarios. However, psychological and neuroscientific studies have shown that emotions exhibit significant ”short-term continuity” characteristics (Davidson 1998; Houben, Van Den Noortgate, and Kuppens 2015). In other words, human emotions are relatively stable over certain periods, with sudden changes being rare. As illustrated in Fig. 1(a), a high cosine similarity is maintained between an anchor sample $x_{i}$ and its neighboring sample $x_{j}$, and even a distant sample $x_{k}$ separated by extended periods (e.g., 60 seconds). Given this nature of emotions, we propose that the definition of positive pairs in CL-based EEG analysis should extend beyond just the anchor and its augmented views. Instead, it should include a broader range of similar samples, especially those that are temporally proximal, as shown in Fig. 1(b) (SCMM). However, existing methods like JCFA (Liu et al. 2024a), which follow the traditional CL paradigm (Chen et al. 2020), may incorrectly pull apart similar but not identical samples, thus failing to capture the inherent correlations of EEG signals.
@@
To address the aforementioned three critical issues, we propose a novel Soft Contrastive Masked Modeling (SCMM) framework for cross-corpus EEG-based emotion recognition. Unlike traditional hard CL shown in Fig. 1(b), SCMM considers emotional continuity and incorporates soft assignments of sample pairs. This approach enables the model to identify the fine-grained relationships between samples in a self-supervised manner, thereby enhancing the generalizability of EEG representations. Comprehensive experiments on three well-recognized datasets demonstrate that SCMM consistently achieves SOTA performance, highlighting its superior capability and stability. In summary, the main contributions of SCMM are outlined as follows:
@@
  • Inspired by the nature of emotions, we propose a novel SCMM framework to address cross-corpus EEG-based emotion recognition. This approach assigns soft weights to sample pairs during contrastive learning to capture the similarity relationships between different samples. As a result, better feature representations of EEG signals are learned in a self-supervised manner. To the best of our knowledge, this is the first study to introduce soft contrastive learning into EEG-based emotion recognition.
@@
  • We also develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships, which is essential for enhancing contrastive learning. Additionally, we introduce an aggregator that weightedly aggregates complementary information from the embeddings of multiple close samples, enabling fine-grained feature learning and improving the model’s overall capability.
@@
  • We conduct extensive experiments on three well-known datasets (SEED, SEED-IV, and DEAP), showing that SCMM achieves consistent SOTA performance compared to eight competitive methods, surpassing the second-best method by an average accuracy of 4.26%.
Methodology
Figure 2: The overall framework of SCMM. The pre-training process of SCMM involves three modules: (1) hybrid masking, (2) soft contrastive learning, and (3) aggregate reconstruction.
Problem Definition
@@
Given an unlabeled pre-training EEG emotion dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$ with $N$ samples, where each sample $x_{i}\in\mathbb{R}^{C\times F}$ contains $C$ channels and $F$-dimensional features, the goal is to learn a nonlinear embedding function $f_{\theta}$. This function is designed to map $x_{i}$ to its representation $h_{i}$ that best describes itself by leveraging the emotional continuity inherent in EEG signals. Ultimately, the pre-trained model is capable of producing generalizable EEG representations that can be effectively used across different EEG emotion datasets.
Model Architecture
@@
The overall framework of SCMM is illustrated in Fig. 2, which includes three main modules: hybrid masking, soft contrastive learning, and aggregate reconstruction. Below, we detail the specific design of each module and the pre-training process of SCMM.
Hybrid Masking
@@
The selection of masking strategies is crucial for CL and masked modeling (Zhang et al. 2024; Liu et al. 2024b). For an input EEG sample $x_{i}\in\mathcal{X}$, most existing methods use random masking (Zhang, Liu, and Zhong 2022) or channel masking (Li et al. 2022) to generate the masked sample $\widetilde{x}_{i}$. The random masking strategy masks samples along the feature dimension, ignoring the inter-channel relationships of EEG signals. While a large masking ratio (e.g., 75%) can mask entire portions of certain channels, it complicates the modeling process due to significant information loss. Conversely, the channel masking strategy masks features across all dimensions of the selected channels, losing the relationships between different dimensional features. Neither approach captures both channel and feature relationships simultaneously. Therefore, we develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships.
@@
Specifically, we first generate a random masking matrix ${\rm Mask}_{R}\in\{0,1\}$ with dimensions $C\times F$, and a channel masking matrix ${\rm Mask}_{C}\in\{0,1\}$ with dimensions $C\times F$, both derived from binomial distributions with masking ratios $r\in(0,1)$. Here, the element values in each row of ${\rm Mask}_{C}$ are either all 1s or all 0s. Next, we generate a probability matrix $U\in[0,1]$ with dimensions $C\times 1$ for hybrid masking, which is drawn from a uniform distribution. The hybrid masking process is defined as:
$\widetilde{x}_{i,c}=\begin{cases}x_{i,c}\odot{\rm Mask}_{R,c}&\text{if }\mu<U_{c}\leq 1\\ x_{i,c}\odot{\rm Mask}_{C,c}&\text{if }0\leq U_{c}\leq\mu\end{cases},$ (1)
@@
where $x_{i,c}$ represents the $c$-th channel of $x_{i}$, and $\widetilde{x}_{i,c}$ is the corresponding masked sample. $\odot$ denotes element-wise multiplication. $U_{c}$ is the probability value in the $c$-th row, and $\mu$ is a probability threshold that controls the weights of the two masking strategies. By integrating the hybrid masking strategy in SCMM, we enhance the diversity of masked samples, encouraging the model to learn richer and more robust feature representations that account for both channel and feature relationships within EEG signals. Figure 3 illustrates the differences between three masking strategies.
Figure 3: Illustrations of different masking strategies. (a) compares various masking strategies, and (b) presents examples of generated masked samples using three strategies. The masking ratio and threshold are set to $r=0.5$ and $\mu=0.1$, respectively.
Soft Contrastive Learning

### v2

@@ Abstract.
Emotion recognition using electroencephalography (EEG) signals has attracted increasing attention in recent years. However, existing methods often lack generalization in cross-corpus settings, where a model trained on one dataset is directly applied to another without retraining, due to differences in data distribution and recording conditions. To tackle the challenge of cross-corpus EEG-based emotion recognition, we propose a novel framework termed Soft Contrastive Masked Modeling (SCMM). Grounded in the theory of emotional continuity, SCMM integrates soft contrastive learning with a hybrid masking strategy to effectively capture emotion dynamics (refer to short-term continuity). Specifically, in the self-supervised learning stage, we propose a soft weighting mechanism that assigns similarity scores to sample pairs, enabling fine-grained modeling of emotional transitions and capturing the temporal continuity of human emotions. To further enhance representation learning, we design a similarity-aware aggregator that fuses complementary information from semantically related samples based on pairwise similarities, thereby improving feature expressiveness and reconstruction quality. This dual design contributes to a more discriminative and transferable representation, which is crucial for robust cross-corpus generalization. Extensive experiments on the SEED, SEED-IV, and DEAP datasets show that SCMM achieves state-of-the-art (SOTA) performance, outperforming the second-best method by an average accuracy of 4.26% under both same-class and different-class cross-corpus settings. The source code is available at https://github.com/Kyler-RL/SCMM.
EEG; Emotion Recognition; Soft Contrastive Learning; Masked Modeling; Cross-Corpus
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755798ccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Emotions are human attitudinal experiences and behavioral responses to objective things, closely related to an individual’s health conditions and behavioral patterns (Wang et al., 2024). Compared to speech (Singh and Goel, 2022), gestures (Noroozi et al., 2018), and facial expressions (Canal et al., 2022), electroencephalography (EEG) offers a more direct and objective measurement of human emotions by capturing brain activity across various scalp locations (Hu et al., 2019). Therefore, researchers have increasingly emphasized EEG-based emotion recognition in recent years (Zhong et al., 2020; Zhao et al., 2021; Zhang et al., 2022, 2023), aiming to advance the development of affective brain-computer interfaces (aBCIs). However, three critical challenges remain to be addressed in current approaches.
@@
Figure 1. (a) An illustration of emotional continuity. We take the sample $x_{i}$ at second $t$ within an EEG trial as the anchor, and calculate the cosine similarity between $x_{i}$ and its neighboring sample $x_{j}$, as well as the distant sample $x_{k}$. High cosine similarities indicate that human emotions remain relatively stable and similar over a certain period. (b) Hard CL vs. SCMM. Traditional hard CL assigns hard values (1 and 0) to positives and negatives when computing the contrastive loss. In contrast, our model generates soft assignments for different sample pairs, taking into account the emotional continuity.
[…]
(a) An illustration of emotional continuity. (b) Hard CL vs. SCMM.
(1) Insufficient Generalization Capability. Most existing EEG-based emotion recognition methods are typically designed for a single dataset, necessitating model retraining when the dataset changes. This requirement significantly limits the scalability and generalizability of the model, hindering its application on different datasets. To address this issue, the concept of cross-corpus has been proposed, which is designed to be generalized across multiple datasets. A cross-corpus model is trained on one dataset and can be directly applied to another without retraining from scratch. This concept, which originated in natural language processing (Schuller et al., 2010; Zhang et al., 2011), has been extended to various domains in recent years (Rayatdoost and Soleymani, 2018; Chien et al., 2020; Ryumina et al., 2022). Although existing EEG-based emotion recognition methods, such as BiDANN (Li et al., 2018b), TANN (Li et al., 2021), and PR-PL (Zhou et al., 2023), have demonstrated superior performance in within-subject or cross-subject tasks within a single dataset, their effectiveness significantly degrades in cross-corpus scenarios, where the differences in data distribution across datasets far exceed the intra-dataset variability (Rayatdoost and Soleymani, 2018).
(2) Modeling Strategy Limitation. Recently, researchers have explored domain adaptation techniques to address cross-corpus EEG-based emotion recognition (He et al., 2022b; Zhou et al., 2025). This is motivated by their efficacy to solve the problem of domain shift (Ganin et al., 2016). Despite initial success, such approaches often require prior access to all labeled source data and unlabeled target data for model training. Considering the difficulties in collecting EEG signals as well as the time and expertise required to label them, the modeling strategy limitation of domain adaptation techniques poses a significant challenge to real-world aBCI applications.
@@
(3) Ignorance of Emotional Continuity. Unlike domain adaptation techniques, contrastive learning (CL) achieves superior performance without relying on labeled data, and has shown great potential in various fields (Chen et al., 2020; Radford et al., 2021; Eldele et al., 2021). Current CL-based methods for EEG-based emotion recognition, such as CLISA (Shen et al., 2022) and JCFA (Liu et al., 2024), consider an anchor and its augmented views as positive pairs, while treating all other samples as negatives. When computing the contrastive loss, the weights for positives and negatives are set to 1 and 0, respectively, as shown in Fig. 1(b) (Hard CL). However, psychological and neuroscientific studies suggest that emotion analysis using brain signals should account for dynamic changes (Davidson, 1998; Houben et al., 2015). Specifically, emotions exhibit significant ”short-term continuity” characteristics, meaning that human emotions are relatively stable over certain periods, with sudden changes being rare. As illustrated in Fig. 1(a), a high cosine similarity is maintained between an anchor sample $x_{i}$ and its neighboring sample $x_{j}$, and even a distant sample $x_{k}$ separated by extended periods (e.g., 60 seconds). Given this nature of emotions, we propose that the definition of positive pairs in CL-based EEG emotion analysis should extend beyond just the anchor and its augmented views. Instead, it should include a broader range of similar samples, especially those that are temporally proximal, as shown in Fig. 1(b) (SCMM). In contrast, existing methods following the traditional CL paradigm (Chen et al., 2020) may incorrectly pull apart similar but not identical samples, thus failing to capture the emotional continuity inherent in EEG signals.
@@
To tackle the aforementioned three critical issues, we propose a novel Soft Contrastive Masked Modeling (SCMM) framework for cross-corpus EEG-based emotion recognition. Unlike traditional hard CL shown in Fig. 1(b), SCMM considers emotional continuity and incorporates soft assignments of sample pairs. This approach enables the model to identify the fine-grained relationships between different samples in a self-supervised manner, thereby enhancing the generalizability of EEG representations. Comprehensive experiments on three well-recognized datasets show that SCMM consistently achieves state-of-the-art (SOTA) performance, demonstrating its superior capability and stability. In summary, the main contributions of SCMM are outlined as follows:
  • We propose a novel SCMM framework to address three key challenges (insufficient generalization capability, modeling strategy limitation, and ignorance of emotional continuity) in cross-corpus EEG-based emotion recognition.
@@
  • Inspired by the nature of emotions, we introduce a soft weighting mechanism that assigns similarity scores to sample pairs to capture the similarity relationships between different samples. As a result, better feature representations of EEG signals are learned in a self-supervised manner.
@@
  • We develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships, which is essential for enhancing contrastive learning. In addition, we introduce a similarity-aware aggregator to fuse complementary information from semantically related samples, enabling fine-grained feature learning and improving the model’s overall capability.
@@
  • We conduct extensive experiments on three well-known datasets (SEED, SEED-IV, and DEAP), demonstrating that SCMM achieves SOTA performance against 10 baselines, with an average accuracy improvement of 4.26% under both same-class and different-class cross-corpus settings.
2. Related Work
2.1. EEG-Based Emotion Recognition
Current approaches for EEG-based emotion recognition mainly rely on two types of experimental protocols: (1) subject-dependent and (2) subject-independent.
(1) The subject-dependent protocol trains and tests models using EEG data from the same subject within a single dataset. For example, Duan et al. (Duan et al., 2013) extracted various emotion-related features from EEG signals and used support vector machine (SVM) and k-nearest neighbors (KNN) for emotion recognition. Similarly, Alsolamy et al. (Alsolamy and Fattouh, 2016) inputted power spectral density (PSD) features into an SVM classifier to predict emotions while listening to the Quran. In terms of deep learning models, Zheng et al. (Zheng and Lu, 2015) trained a deep belief network (DBN) using differential entropy (DE) features extracted from multi-channel EEG signals for subject-dependent emotion classification. Additionally, Song et al. (Song et al., 2018) proposed a dynamical graph convolutional network (DGCNN) that dynamically learns the intrinsic relationship between different EEG channels to enhance the model’s discriminative ability. In general, the subject-dependent protocol tends to achieve superior performance due to its potential to introduce information leakage. However, this protocol fails to take into account the significant individual differences of EEG signals, thus limiting its practical applications.
(2) The subject-independent protocol trains and tests models using EEG data from different subjects within a single dataset. Since transfer learning has demonstrated its potential in addressing the problem of domain shift, a series of methods have adopted it for subject-independent EEG-based emotion recognition. For example, Li et al. (Li et al., 2018b) proposed a bi-hemispheres domain adversarial neural network (BiDANN) that considers distribution shift between training and testing data and cerebral hemispheres. Following this, a novel transferable attention neural network (TANN) (Li et al., 2021) was introduced to learn the emotional discriminative information of EEG signals. To simultaneously adapt the marginal distribution and the conditional distribution, Li et al. (Li et al., 2019) proposed a joint distribution adaptation network (JDA) for subject-independent EEG-based emotion recognition. Similarly, Chen et al. (Chen et al., 2021) introduced a multi-source marginal distribution adaptation network (MS-MDA) to capture both domain-invariant and domain-specific features of emotional EEG signals. Further, Zhou et al. (Zhou et al., 2023) proposed a novel prototypical representation-based pairwise learning framework (PR-PL) to address individual differences and noisy labeling in emotional EEG signals. Despite the progress made in dealing with individual differences, these methods struggle to mitigate the distributional differences across datasets, resulting in significant performance degradation in cross-corpus scenarios.
2.2. Cross-Corpus EEG Emotion Recognition
To alleviate the large distributional differences across datasets, researchers have attempted to use domain adaptation techniques for cross-corpus EEG-based emotion recognition. For example, He et al. (He et al., 2022b) proposed an adversarial discriminative temporal convolutional network (AD-TCN) that integrates the adversarial discriminative learning into a temporal convolutional network for enhancing distribution matching. Meanwhile, Li et al. (Li et al., 2024) proposed a novel distillation-based domain generalization network (DBDG) to learn the discriminative and generalizable emotional features. Moreover, Zhou et al. (Zhou et al., 2025) introduced an EEG-based emotion style transfer network (E2STN) that contains the content information of the source domain and the style information of the target domain, achieving superior performance in cross-corpus scenarios. However, these methods require prior access to all labeled source data and unlabeled target data for model training, which is not feasible in practical applications due to the difficulties in collecting and labeling EEG data. Therefore, recent studies have sought to leverage contrastive learning for cross-corpus EEG-based emotion recognition. A noteworthy attempt is JCFA (Liu et al., 2024), which performs joint contrastive learning across three domains to align the time- and frequency-based embeddings of the same EEG sample in the latent time-frequency space, achieving SOTA performance in cross-corpus EEG-based emotion recognition tasks. However, such approaches following the traditional CL paradigm (Chen et al., 2020) fail to capture the emotional continuity inherent in EEG signals, resulting in relatively limited performance in cross-corpus scenarios. Therefore, this study aims to develop a sufficiently generalized model without relying on labeled data, which can be directly applied to different EEG emotion datasets without retraining from scratch, and achieve accurate and efficient cross-corpus generalization.
3. Problem Formulation
@@
Given an unlabeled pre-training EEG emotion dataset $X=\{x_{i}\}_{i=1}^{N}$ with $N$ samples, where each sample $x_{i}\in\mathbb{R}^{C\times F}$ contains $C$ channels and $F$-dimensional features, the goal is to learn a nonlinear embedding function $f_{\theta}$. This function is designed to map $x_{i}$ to its representation $h_{i}$ that best describes itself by leveraging the emotional continuity inherent in EEG signals. Ultimately, the pre-trained model is capable of producing generalizable EEG representations that can be effectively used across different datasets.
@@
Figure 2. The overall framework of SCMM. The pre-training process of SCMM involves three modules: (1) hybrid masking, (2) soft contrastive learning, and (3) aggregate reconstruction. Best viewed in color.
[…]
The overall framework of SCMM.
4. Methodology
The overall framework of SCMM is illustrated in Fig. 2, which includes three main modules: hybrid masking, soft contrastive learning, and aggregate reconstruction. Below, we will detail the specific design of each module and the pre-training process of SCMM.
4.1. Hybrid Masking
@@
The selection of masking strategies is crucial for CL and masked modeling. For an input EEG sample $x_{i}\in X$, existing methods use random masking (Zhang et al., 2022) or channel masking (Li et al., 2022b) to generate the masked sample $\widetilde{x}_{i}$. The random masking strategy masks samples along the feature dimension, ignoring the inter-channel relationships of multi-channel EEG signals. While a large masking ratio (e.g., 75%) can mask entire portions of certain channels, it complicates the modeling process due to significant information loss. Conversely, the channel masking strategy masks features across all dimensions of the selected channels, losing the relationships between different dimensional features. Neither approach captures both channel and feature relationships simultaneously. Therefore, we develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships.
@@
Specifically, we first generate a random masking matrix ${\rm Mask}_{R}\in\{0,1\}$ with dimensions $C\times F$ and a channel masking matrix ${\rm Mask}_{C}\in\{0,1\}$ with dimensions $C\times F$, both derived from binomial distributions with the same masking ratio $r\in(0,1)$. Here, the element values in each row of ${\rm Mask}_{C}$ are either all 1s or all 0s. Next, we generate a probability matrix $U\in[0,1]$ with dimensions $C\times 1$ for hybrid masking, which is drawn from a uniform distribution. The hybrid masking process is defined as:
(1) $\widetilde{x}_{i,c}=\begin{cases}x_{i,c}\odot{\rm Mask}_{R,c}&\text{if }\mu<U_{c}\leq 1\\ x_{i,c}\odot{\rm Mask}_{C,c}&\text{if }0\leq U_{c}\leq\mu\end{cases},$
@@
where $\odot$ denotes element-wise multiplication. $x_{i,c}$ represents the DE features of the $c$-th channel of $x_{i}$, and $\widetilde{x}_{i,c}$ is the corresponding masked sample. $U_{c}$ represents the probability value of $U$ in the $c$-th row, and $\mu$ is a probability threshold that controls the weights of the two masking strategies. By integrating the hybrid masking strategy in SCMM, we enhance the diversity of masked samples, enabling the model to learn richer feature representations that account for both channel and feature relationships of EEG signals. Figure 3 illustrates the differences between the three masking strategies. More details are presented in Appendix D.
Figure 3. Comparison of different masking strategies. The masking ratio and threshold are set to $r=0.5$ and $\mu=0.1$, respectively. Best viewed in color. Zoom in for a better view.
[…]
Comparison of different masking strategies.
4.2. Soft Contrastive Learning